# Removing singularities for fully nonlinear PDEs

### v1

@@ 1.1 Monge-Ampère
This result originally follows from the regularity theory of Monge-Ampère. The singular set is well known by Pogorelov’s Hessian estimate [P64, P78] to be the set of points where $u$ fails strict convexity, and is closed in $B_{1}(0)$. It follows from Caffarelli [C90] that $u$ only equals a linear function on sets with no extremal points: subspaces. For example, the Pogorelov [P78] singular solution $u(x)=|(x_{1},x_{2})|^{1+\alpha}f(x_{3})$ is singular along a full line. The dimension of degeneration is less than $n/2$, as follows from Caffarelli [C93], where examples of such dimensions were also produced. It follows from Mooney [M15] that the collection of $k$-dimensional degenerate subspaces has Hausdorff dimension less than $n-k$, and that $\text{sing}(u)$ has sharp upper bound of dimension $n-1$. In dimension two, Alexandrov showed strict convexity [A42].
@@ 1.2 General PDEs: minimal surface, special Lagrangian equation
$|D^{k}u|\leq C(k,B,\Omega^{\prime},\|u\|_{L^{\infty}(B_{1}(0))}),\qquad k=0,1,2,3\dots{}$ UNKNOWN (1.6)
@@
There are many other nonlinear PDEs which have Jacobi inequalities, at least under some a priori conditions on the solutions. To list a few recent ones, there are the Lagrangian mean curvature equations [B21, L23a, Z23, BW24a, BW24b], generalized special Lagrangian equation [Z24], the Hessian quotient equation [L23b, L24], scalar curvature equations [CJZ24, QZ24, W24], and 2-Hessian equation with dynamic semi-convexity [ShY23].
@@ 1.4 Discussion
An earlier, standard result is the removability of point singularities for $C^{\infty}$-solvable equations, i.e. those PDEs which have classical solutions to Dirichlet problems with smooth boundary data, as in [CNS85]. A viscosity solution has a point singularity at $p\in\Omega$ if $u\in C^{\infty}(\overline{\Omega}/\setminus\{p\})\cap C^{0}(\overline{\Omega})$. The result states that if a $C^{\infty}$-solvable equation has a solution with a point singularity, then it is smooth. The proof is by the comparison principle.
@@ 1.5 Outline of the doubling proof
Our approach follows [SY23] and compares with a “1D Green’s function” $(1/x_{1}-1)_{+}$. In [SY23], the Green’s function vanishes on the boundary, so the small defective set is near the pole of the Green’s function. But in our situation, the Green’s function is nonvanishing near $x_{2}=\pm 1$ as well as the pole $x_{1}=0$. The doubling inequality only controls $b(D^{2}u,Du)$ in terms of its values on three sides of a box. But provided the singular set, such as a half-line, satisfies the single side condition, it follows that $b(D^{2}u,Du)$ is bounded on the three bad sides, and we obtain a global estimate.

### v2

@@ 1.1 Monge-Ampère
This result originally follows from the regularity theory of Monge-Ampère. The singular set is well known by Pogorelov’s Hessian estimate [P64, P78] to be the set of points where $u$ fails strict convexity, and is closed in $B_{1}(0)$. It follows from Caffarelli [C90] that $u$ only equals a linear function on sets with no extremal points: subspaces. For example, the Pogorelov [P78] singular solution $u(x)=|(x_{1},x_{2})|^{1+\alpha}f(x_{3})$ is singular along a full line. The dimension of degeneration is less than $n/2$, as follows from Caffarelli [C93], where examples of such dimensions were also produced. It follows from Mooney [M15] that the collection of $k$-dimensional degenerate subspaces has Hausdorff dimension less than $n-k$, and that $\text{sing}(u)$ has sharp upper bound of dimension $n-1$. In dimension two, Alexandrov showed strict convexity [A42]. Singular set propagation to the boundary was obtained for Gaussian curvature and curvature functions by Smith [S12, S24]. Namely, the singular set is the union of the convex hull of closed subsets of the domain boundary.
@@ 1.2 General PDEs: minimal surface, special Lagrangian equation
$|D^{k}u|\leq C(k,B,\Omega^{\prime},\Omega,\|u\|_{L^{\infty}(B_{1}(0))}),\qquad k=0,1,2,3\dots{}$ UNKNOWN (1.6)
@@
There are many other nonlinear PDEs which have Jacobi or Jacobi type inequalities, at least under some a priori conditions on the solutions. To list a few recent ones, there are the Lagrangian mean curvature equations [B21, L23a, Z23, BW24a, BW24b], generalized special Lagrangian equation [Z24], the Hessian quotient equation [L23b, L24], scalar curvature equations [CJZ24, QZ24, W24], and 2-Hessian equation with dynamic semi-convexity [ShY23].
@@ 1.4 Discussion
An earlier, standard result is the removability of point singularities for $C^{\infty}$-solvable equations, i.e. those PDEs which have classical solutions to Dirichlet problems with smooth boundary data, such as concave PDEs as in [CNS85]. A viscosity solution has a point singularity at $p\in\Omega$ if $u\in C^{\infty}(\overline{\Omega}/\setminus\{p\})\cap C^{0}(\overline{\Omega})$. The result states that if a $C^{\infty}$-solvable equation has a solution with a point singularity, then it is smooth. The proof is by the comparison principle.
@@ 1.5 Outline of the doubling proof
Our approach follows [SY23] and compares with a “1D Green’s function” $(1/|x_{1}|-1)_{+}$. In [SY23], the Green’s function vanishes on the boundary, so the defective set is near the pole of the Green’s function. But in our situation, the Green’s function is nonvanishing near $x_{2}=\pm 1$ as well as the pole $x_{1}=0$. The doubling inequality only controls $b(D^{2}u,Du)$ in terms of its values on these sides of a box. But provided the singular set, such as a half-line, satisfies the single side condition, it follows that $b(D^{2}u,Du)$ is bounded on the three bad sides, and we obtain a global estimate.